l. At an amusement park there is a ride in which cylindrically shaped chambers spin around a central axis. People sit in seats facing the axis, their backs against the outer wall. At one instant the outer wall moves at a speed of 3.37 m/s, and a 59.1-kg person feels a 436-N force pressing against his back. What is the radius of a chamber?
Answer #1

Solution:

Step 1: Given:

Speed(v)3.37m/s, mass(m)-59.1kg, Force 436N

Radius (r)= ??

Using equation: F= \frac{mv^2}{r}

so, r= \frac{mv^2}{F}

Putting values in above equation we get:

r= \frac{59.1*(3.37)^2}{436}

r= 1.54 m